# Transolver++: An Accurate Neural Solver for PDEs on Million-Scale Geometries

### v1

@@ Results
We also observe that GNNs often degenerate quickly when handling large-scale meshes due to their geometric instability (instabilityGNN). Additionally, Geo-FNO also performs poorly across nearly all datasets, as it attempts to map irregular meshes to uniform latent grids, a task that becomes especially difficult when the number of meshes scales up. These failures of our baselines further highlight the challenge of physics learning on million-scale geometries, while Transolver++ provides a practical solution, advancing an essential step to industrial applications.
@@ 5.3 Model Analysis
From the car mesh comparison in Figure 1(b), we can observe that large geometries differ significantly from smaller-scale ones in terms of details, which directly affects the accuracy of the simulation. While previous models (li2021fourier) claim to be resolution-invariant and aim to apply directly to large-scale geometries, our experiments in Figure 6 (c.2) show that without training on large meshes, the model will fall short in the fine-grained physics modeling, resulting in an insufficient and limited performance, which further necessitates the capability of handling larger geometries.
@@ 6 Conclusion
Impact Statement
This paper aims at advancing neural PDE solvers to industrial applications. Building on the previous advancement of Transolver, we present Transolver++ as an accurate neural solver for PDEs on million-scale geometries, which achieves consistent state-of-the-art in both standard benchmarks and industrial-level tasks with significant promotion. Also, Transolver++ pushes the single-GPU input size capacity to the million scale for the first time and achieves linear scalability with resources, showing favorable practicability to industrial manufacturing. Our work only focuses on the scientific problem, so we believe there is no potential ethical risk.
@@ Appendix A Full Results on Standard Benchmarks
For models marked with an asterisk (*), we carefully reproduced the results by running the models more than three times and enabled fair and reliable comparison by maintaining the model parameter counts within a closely comparable range with other models. Specifically, we adjust LNO and reduce its default dimension from 256 to 192 in elasticity, as its parameter count is already three times larger than the other largest models. As discussed in previous studies (wu2024Transolver), Transformer-based models usually benefit from large parameter sizes. This approach minimizes the potential influence of parameter variations and highlights the performance improvements achieved by our method under similar computation costs.
@@
FNO (li2021fourier) / / / / 0.1556 0.0108
U-FNO (Wen2021UFNOA) 0.0239 0.0039 0.0269 0.0056 0.2231 0.0183
geo-FNO (Li2022FourierNO) 0.0229 0.0074 0.0138 0.0067 0.1556 0.0108
U-NO (rahman2022u) 0.0258 0.0034 0.0078 0.0100 0.1713 0.0113
F-FNO (anonymous2023factorized) 0.0263 0.0047 0.0078 0.0070 0.2322 0.0077
LSM (wu2023LSM) 0.0218 0.0025 0.0059 0.0050 0.1535 0.0065
LNO* (wang2024LNO) 0.0069 0.0029 0.0053 0.0031 0.0830 0.0063
Galerkin (Cao2021ChooseAT) 0.0240 0.0120 0.0118 0.0098 0.1401 0.0084
HT-Net (anonymous2023htnet) / 0.0333 0.0065 0.0059 0.1847 0.0079
OFormer (li2023transformer) 0.0183 0.0017 0.0183 0.0168 0.1705 0.0124
GNOT (hao2023gnot) 0.0086 0.0336 0.0076 0.0047 0.1380 0.0105
FactFormer (li2023scalable) / 0.0312 0.0071 0.0060 0.1214 0.0109
ONO (anonymous2023improved) 0.0118 0.0048 0.0061 0.0052 0.1195 0.0076
Transolver (wu2024Transolver) 0.0064 0.0013 0.0053 0.0033 0.0900 0.0058
@@ B.1 Benchmarks
This benchmark is generated by the simulations of the stress field in a hyper-elastic solid body under tension, which is governed by a stress-strain relationship using the Rivlin-Saunders material model (Li2022FourierNO). Each case involves a unit cell of 972 points with a void in the middle , clamped at the bottom edge and subjected to tension on the top. Elasticity contains a total of 1200 samples, with 1000 for training and 200 for testing.
@@
This benchmark is generated by the simulation of a plastic forging problem, where a block is impacted by a frictionless die moving at a constant speed (Li2022FourierNO). An elastoplastic constitutive model is adopted to model this physical system with 900 training and 80 testing samples. The input is the external force on every mesh with the shape of 101 $\times$ 31, and the output is the time-dependent deformation and mesh grid over 20 timesteps.
@@
This benchmark is generated from simulations of transonic flow over an airfoil, governed by Euler’s equations (Li2022FourierNO). The whole field is discretized to unstructured meshes in the shape of 221 $\times$ 51 as the input and the output is the corresponding Mach number on these meshes. The dataset includes 1000 training samples and 200 test samples that are based on the initial NACA-0012 shape.
@@
This benchmark consists of simulations of incompressible flow in a pipe, governed by the Navier-Stokes equation with viscosity $\nu=0.005$ (Li2022FourierNO). The pipe has a length of 10 and a width of 1, with its centerline parameterized by cubic polynomials determined by five control nodes. The dataset also contains 1000 training and 200 testing samples, whose inputs are the mesh point locations 129 $\times$ 129 and outputs are the horizontal velocity field.
@@
This benchmark consists of simulations of the 2D Navier-Stokes equations in vorticity form on the unit torus $(0,1)^{2}$ (li2021fourier). The objective is, given the past velocity, to predict future velocity for 10 steps on discretized meshes in the shape of 64 $\times$ 64. The inputs are the velocity for the past 10 steps, while the outputs provide the future velocity for 10 timesteps. The dataset also includes 1000 samples for training and 200 for testing.
@@
This benchmark consists of simulations of the steady-state Darcy Flow in two dimensions, governed by a second-order elliptic equation on the unit square (Li2022FourierNO). The input is the structure of the porous medium and the output is the corresponding fluid pressure. The dataset contains 1000 training samples and 200 testing samples, generated using a second-order finite difference scheme on 421 $\times$ 421 uniform grids and later downsampled to $85\times 85$.
@@
This benchmark (elrefaie2024drivaernet++) is a large and comprehensive dataset for aerodynamic car design, featuring high-fidelity computational fluid dynamics (CFD) simulations. It includes over 8,000 car designs with various configurations of wheel and underbody design. To ensure efficiency while maintaining diversity, we select 200 representative cases from these designs. Notably, in our experiments, DrivAerNet++ is then divided into two subsets with different levels of resolution. The first subset, as shown in Table 6, only consists of surface meshes, where each mesh point is characterized by its 3D position $(x,y,z)$, surface normal vector $(u_{x},u_{y},u_{z})$, and signed distance function (SDF). The output for this subset is the surface pressure on these meshes. The second subset provides a full 3D pressure and velocity field, significantly increasing the dataset’s complexity, with the number of mesh points reaching approximately 2.5 million and requiring the model to predict surface pressure, velocity and pressure of the surrounding area. The dataset consists of 190 training samples and 10 test samples, offering a rigorous benchmark for evaluating aerodynamic modeling at different levels of fidelity.
@@ B.2.2 Large-Scale Datasets
Table 7: Implementation details of Transolver++ including training and model configuration. Training configurations are identical to previous methods (wu2024Transolver; hao2023gnot; anonymous2023geometryguided; elrefaie2024drivaernet++) and shared in all baselines. $\mathcal{L}_{\mathrm{v}}$ and $\mathcal{L}_{\mathrm{s}}$ refer to the loss on volume (surrounding area) and surface physics fields respectively.
@@
DrivAerNet++ Surf $\mathcal{L}_{\mathrm{s}}$ (DBLP:journals/corr/KingmaB14)
@@ B.3.1 Standard Benchmarks
We try to be as loyal to the original settings of baselines as possible. However, when the model parameters count is too large to conduct a fair comparison, we would change either the number of blocks or the hidden dimension to ensure a fair comparison. As we mentioned before, Transformer-based models usually benefit from large parameter sizes (wu2024Transolver). Thus, the alignment of parameter size is essential to control the variable and highlight the performance difference caused by architecture design. Here are detailed implementations.
@@
Besides, we carefully review the original papers of these baseline models to make sure that they are fully tuned on their hyper-parameters. For all the models whose data settings align exactly with ours, such as FNO (li2021fourier), Geo-FNO (Li2022FourierNO), GNOT (hao2023gnot), OFormer (li2023transformer), Transolver (wu2024Transolver), we directly adopt the results reported in their respective publications. For other models, we refer to the results presented in LSM (wu2023LSM) and Transolver (wu2024Transolver), as these works have conducted comprehensive and rigorous hyper-parameter tuning to ensure a fair and reliable comparison.
@@ B.3.2 Industrial Applications
For DrivAerNet++ and AirCraft datasets, most of the baseline models are unable to handle million-scale meshes directly. To address this, we subsample the dataset to 50k meshes point and reconstruct the meshes using K-Nearest Neighbors (KNN) (peterson2009knearest) method to preserve essential geometric relationships.
@@
In terms of model configurations, each model has been extensively tuned on their hyperparameters, especially graph neural networks (GNNs) for their extreme instability (instabilityGNN) when dealing with million-scale meshes. To minimize human bias, we employ a standardized tuning strategy by systematically changing the hidden dimension from {64, 128, 256, 512}, the number of layers from {2, 4, 6, 8, 10} along with tuning model-specific hyperparameters. During this process, we observe that almost all GNNs experienced geometric instability (morris2023geometric) when applied to the DrivAerNet++ Full dataset with approximately 2.5 million mesh points per case. Out of all baselines, Transolver achieves the second-best performance, benefiting from its Slice-Deslice mechanism, which improves its stability and scalability, while Transolver++ surpasses it by introducing eidetic physical states and a parallelism framework that can directly handle million-scale meshes with enhanced efficiency. In Transolver++, we set the number of slices to 32 and its channels to 256 with 4 layers of Transolver++ block to balance efficiency and performance. And we only need to adopt parallelism to Transolver++ for the DrivAerNet++ Full benchamrk on only 4 A100 GPUs at most.

### v2

@@ Results
We also observe that GNNs often degenerate quickly when handling large-scale meshes due to their geometric instability (Klabunde & Lemmerich, 2023). Additionally, Geo-FNO also performs poorly across nearly all datasets, as it attempts to map irregular meshes to uniform latent grids, a task that becomes especially difficult when the number of meshes scales up. These failures of our baselines further highlight the challenge of physics learning on million-scale geometries, while Transolver++ provides a practical solution, advancing an essential step to industrial applications.
@@ 5.3 Model Analysis
From the car mesh comparison in Figure 1(b), we can observe that large geometries differ significantly from smaller-scale ones in terms of details, which directly affects the accuracy of the simulation. While previous models (Li et al., 2021) claim to be resolution-invariant and aim to apply directly to large-scale geometries, our experiments in Figure 6 (c.2) show that without training on large meshes, the model will fall short in the fine-grained physics modeling, resulting in an insufficient and limited performance, which further necessitates the capability of handling larger geometries.
@@ Appendix A Full Results on Standard Benchmarks
For models marked with an asterisk (*), we carefully reproduced the results by running the models more than three times and enabled fair and reliable comparison by maintaining the model parameter counts within a closely comparable range with other models. Specifically, we adjust LNO and reduce its default dimension from 256 to 192 in elasticity, as its parameter count is already three times larger than the other largest models. As discussed in previous studies (Wu et al., 2024), Transformer-based models usually benefit from large parameter sizes. This approach minimizes the potential influence of parameter variations and highlights the performance improvements achieved by our method under similar computation costs.
@@
FNO (2021) / / / / 0.1556 0.0108
U-FNO (2022) 0.0239 0.0039 0.0269 0.0056 0.2231 0.0183
geo-FNO (2022) 0.0229 0.0074 0.0138 0.0067 0.1556 0.0108
U-NO (2023) 0.0258 0.0034 0.0078 0.0100 0.1713 0.0113
F-FNO (2023) 0.0263 0.0047 0.0078 0.0070 0.2322 0.0077
LSM (2023) 0.0218 0.0025 0.0059 0.0050 0.1535 0.0065
LNO* (2024) 0.0069 0.0029 0.0053 0.0031 0.0830 0.0063
Galerkin (2021) 0.0240 0.0120 0.0118 0.0098 0.1401 0.0084
HT-Net (2022) / 0.0333 0.0065 0.0059 0.1847 0.0079
OFormer (2023c) 0.0183 0.0017 0.0183 0.0168 0.1705 0.0124
GNOT (2023) 0.0086 0.0336 0.0076 0.0047 0.1380 0.0105
FactFormer (2023d) / 0.0312 0.0071 0.0060 0.1214 0.0109
ONO (2024) 0.0118 0.0048 0.0061 0.0052 0.1195 0.0076
Transolver (2024) 0.0064 0.0013 0.0053 0.0033 0.0900 0.0058
@@ B.1 Benchmarks
This benchmark is generated by the simulations of the stress field in a hyper-elastic solid body under tension, which is governed by a stress-strain relationship using the Rivlin-Saunders material model (Li et al., 2022). Each case involves a unit cell of 972 points with a void in the middle , clamped at the bottom edge and subjected to tension on the top. Elasticity contains a total of 1200 samples, with 1000 for training and 200 for testing.
@@
This benchmark is generated by the simulation of a plastic forging problem, where a block is impacted by a frictionless die moving at a constant speed (Li et al., 2022). An elastoplastic constitutive model is adopted to model this physical system with 900 training and 80 testing samples. The input is the external force on every mesh with the shape of 101 $\times$ 31, and the output is the time-dependent deformation and mesh grid over 20 timesteps.
@@
This benchmark is generated from simulations of transonic flow over an airfoil, governed by Euler’s equations (Li et al., 2022). The whole field is discretized to unstructured meshes in the shape of 221 $\times$ 51 as the input and the output is the corresponding Mach number on these meshes. The dataset includes 1000 training samples and 200 test samples that are based on the initial NACA-0012 shape.
@@
This benchmark consists of simulations of incompressible flow in a pipe, governed by the Navier-Stokes equation with viscosity $\nu=0.005$ (Li et al., 2022). The pipe has a length of 10 and a width of 1, with its centerline parameterized by cubic polynomials determined by five control nodes. The dataset also contains 1000 training and 200 testing samples, whose inputs are the mesh point locations 129 $\times$ 129 and outputs are the horizontal velocity field.
@@
This benchmark consists of simulations of the 2D Navier-Stokes equations in vorticity form on the unit torus $(0,1)^{2}$ (Li et al., 2021). The objective is, given the past velocity, to predict future velocity for 10 steps on discretized meshes in the shape of 64 $\times$ 64. The inputs are the velocity for the past 10 steps, while the outputs provide the future velocity for 10 timesteps. The dataset also includes 1000 samples for training and 200 for testing.
@@
This benchmark consists of simulations of the steady-state Darcy Flow in two dimensions, governed by a second-order elliptic equation on the unit square (Li et al., 2022). The input is the structure of the porous medium and the output is the corresponding fluid pressure. The dataset contains 1000 training samples and 200 testing samples, generated using a second-order finite difference scheme on 421 $\times$ 421 uniform grids and later downsampled to $85\times 85$.
@@
This benchmark (Elrefaie et al., 2024) is a large and comprehensive dataset for aerodynamic car design, featuring high-fidelity computational fluid dynamics (CFD) simulations. It includes over 8,000 car designs with various configurations of wheel and underbody design. To ensure efficiency while maintaining diversity, we select 200 representative cases from these designs. Notably, in our experiments, DrivAerNet++ is then divided into two subsets with different levels of resolution. The first subset, as shown in Table 6, only consists of surface meshes, where each mesh point is characterized by its 3D position $(x,y,z)$, surface normal vector $(u_{x},u_{y},u_{z})$, and signed distance function (SDF). The output for this subset is the surface pressure on these meshes. The second subset provides a full 3D pressure and velocity field, significantly increasing the dataset’s complexity, with the number of mesh points reaching approximately 2.5 million and requiring the model to predict surface pressure, velocity and pressure of the surrounding area. The dataset consists of 190 training samples and 10 test samples, offering a rigorous benchmark for evaluating aerodynamic modeling at different levels of fidelity.
@@ B.2.2 Large-Scale Datasets
Table 7: Implementation details of Transolver++ including training and model configuration. Training configurations are identical to previous methods (Wu et al., 2024; Hao et al., 2023; Deng et al., 2024; Elrefaie et al., 2024) and shared in all baselines. $\mathcal{L}_{\mathrm{v}}$ and $\mathcal{L}_{\mathrm{s}}$ refer to the loss on volume (surrounding area) and surface physics fields respectively.
@@
DrivAerNet++ Surf $\mathcal{L}_{\mathrm{s}}$ (2015)
@@ B.3.1 Standard Benchmarks
We try to be as loyal to the original settings of baselines as possible. However, when the model parameters count is too large to conduct a fair comparison, we would change either the number of blocks or the hidden dimension to ensure a fair comparison. As we mentioned before, Transformer-based models usually benefit from large parameter sizes (Wu et al., 2024). Thus, the alignment of parameter size is essential to control the variable and highlight the performance difference caused by architecture design. Here are detailed implementations.
@@
Besides, we carefully review the original papers of these baseline models to make sure that they are fully tuned on their hyper-parameters. For all the models whose data settings align exactly with ours, such as FNO (Li et al., 2021), Geo-FNO (Li et al., 2022), GNOT (Hao et al., 2023), OFormer (Li et al., 2023c), Transolver (Wu et al., 2024), we directly adopt the results reported in their respective publications. For other models, we refer to the results presented in LSM (Wu et al., 2023) and Transolver (Wu et al., 2024), as these works have conducted comprehensive and rigorous hyper-parameter tuning to ensure a fair and reliable comparison.
@@ B.3.2 Industrial Applications
For DrivAerNet++ and AirCraft datasets, most of the baseline models are unable to handle million-scale meshes directly. To address this, we subsample the dataset to 50k meshes point and reconstruct the meshes using K-Nearest Neighbors (KNN) (Peterson, 2009) method to preserve essential geometric relationships.
@@
In terms of model configurations, each model has been extensively tuned on their hyperparameters, especially graph neural networks (GNNs) for their extreme instability (Klabunde & Lemmerich, 2023) when dealing with million-scale meshes. To minimize human bias, we employ a standardized tuning strategy by systematically changing the hidden dimension from {64, 128, 256, 512}, the number of layers from {2, 4, 6, 8, 10} along with tuning model-specific hyperparameters. During this process, we observe that almost all GNNs experienced geometric instability (Morris et al., 2023) when applied to the DrivAerNet++ Full dataset with approximately 2.5 million mesh points per case. Out of all baselines, Transolver achieves the second-best performance, benefiting from its Slice-Deslice mechanism, which improves its stability and scalability, while Transolver++ surpasses it by introducing eidetic physical states and a parallelism framework that can directly handle million-scale meshes with enhanced efficiency. In Transolver++, we set the number of slices to 32 and its channels to 256 with 4 layers of Transolver++ block to balance efficiency and performance. And we only need to adopt parallelism to Transolver++ for the DrivAerNet++ Full benchamrk on only 4 A100 GPUs at most.